OpenStudy (anonymous):

The price per barrel of crude oil in constant 2008 dollars can be approximated by P(t) = 0.45t^2 − 12t + 105 dollars. (0 ≤ t ≤ 28) where t is time in years since the start of 1980.† Find P'(t) and P'(26). What does the answer tell you about the price of crude oil? The price of a barrel of crude oil was increasing at a rate of $ ___ per year in ___

OpenStudy (anonymous):

Well, find the derivative and put in 26.

OpenStudy (anonymous):

P'(t) = 0.9t - 12 P'(26) = 0.9*26 - 12 = 11.4 So (assuming t = 0 in 1980), in 2006 the price of oil was increasing by $11.40.